A local moving service recently purchased a van by securing a loan with semiannual payments of $2700 per semiannual period for 6 years at 13% per year compounded semiannually. What was the purchase price of this van? (Round your answer to the nearest cent.)
$
Answer #1

PV =

P = $2700, n = 6 * 2 = 12 semi-annual periods, r = 13%/2 = 6.5% (semi-annually)

PV = 2700 * 8.1587

PV = $22,028.6